OpenStudy (anonymous):

In a 45°-45°-90° triangle, the length of the hypotenuse is 11. Find the length of one of the legs.

OpenStudy (anonymous):

\[L=\frac{ 1 }{ 2 }H \sqrt{2}\]

OpenStudy (anonymous):

A) \[11 \sqrt{3}\div 2\]

OpenStudy (anonymous):

B) \[11\sqrt{2}\div 2\]

OpenStudy (anonymous):

C) \[11 \sqrt{2}\]

OpenStudy (anonymous):

You don't really need to put the choices out, just plug in the H of the formula i gave you and you'll be good...

OpenStudy (anonymous):

i dont have a calculator on me

OpenStudy (anonymous):

You don't need one...

OpenStudy (anonymous):

im not good with doing numbers in my head

OpenStudy (anonymous):

H=Hypotenuse. So put the H in for the formula and tell me what you get

OpenStudy (anonymous):

i know what H is

OpenStudy (anonymous):

wait so its C?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[L=H \sqrt{2} \div 2\]

OpenStudy (anonymous):

\[H=11\]

OpenStudy (anonymous):

I hope you know that I'm just re-stating my first formula, re-arranged in a different way, right?